OpenStudy (lia22):

WILL MEDAL AND FAN!!!!!! What is the volume of a box whose dimensions are 5'x6'x7' ?

OpenStudy (marziman):

So we're talking about a cube.

OpenStudy (marziman):

Wait no, a rectangular prism.

OpenStudy (marziman):

Do you know how to get the volume?

OpenStudy (lia22):

._.

OpenStudy (marziman):

Volume = W*H*L

OpenStudy (lia22):

Volume of wut?

OpenStudy (lia22):

So w=5 L=7 and H=6 ?

OpenStudy (marziman):

I believe so, is there an image?

OpenStudy (lia22):

No

OpenStudy (marziman):

Well it says it's a box, meaning this is a 3 dimmensional object AND it wants the volume so that's how we know it's a 3D object

OpenStudy (lia22):

Oh okay

OpenStudy (marziman):

We know it's a rectangular prism which is a fancy word for "3D rectangle" or "Rectangle box" because all of the sides are not the same

OpenStudy (lia22):

V=W*L*H V=210

OpenStudy (marziman):

OH DERP I know what they are now

OpenStudy (lia22):

Wut?

OpenStudy (marziman):

The length on a rectangle is the longest side, the width is the second shortest size and the height is the shortest side.

OpenStudy (lia22):

L= 7 W=6 H=5

OpenStudy (marziman):

Yes :D

OpenStudy (marziman):

And if you want to get technical the length * the width is one base of a rectangular prism

OpenStudy (marziman):